Use the method of undetermined coefficients to solve the given nonhomogeneous system. let 3 X + 6 X(t) =
Answer #1

=- c_1 e^(3 t)+ c_2 e^(7 t) + (9 e^t)/4

= 3 c_1 e^(3 t) +9 c_2 e^(7 t) - (21 e^t)/4

x = 13 ༧ག )x+ 1)e 6 4. Eigenvalues of ( .. 9 03) 4- 43 9 6-8/=0. ། 72 -101 + 24 - 3=0 12-107 +2120-1= 3,7. 7.1 = 3 =) Y3 3 *라 14. JU -33-674 ) att = *-7e-2218 e아 18 ( ( 2). 14) Xplt)=AF) -7e-2-12 getl4 e-5 18 -2leH4 3L -cie X (H) = +C2e가 +9e/4 3ce3